OpenStudy (coolio100):

The shortest side of a right triangle measures 7 m. The lengths of the other two sides are consecutive even integers. Find the lengths of the other two sides.

OpenStudy (coolio100):

How do we figure out the second legs size? Or do we have to find the hypotenuse first?

OpenStudy (jiteshmeghwal9):

Solve for x and find the length of the other side

OpenStudy (coolio100):

X=45/4?

OpenStudy (jiteshmeghwal9):

Length of shortest side is 7 m Length of the other side is 2x m Length of hypotenuse become 2x+2 m Apply Pythagoras theorem (7)^2+(2x)^2=(2x+2)^2

OpenStudy (jiteshmeghwal9):

@Coolio100

OpenStudy (coolio100):

O that's why it's not making any sense the shortest size measures 6 not 7! Sorry I miss typed

OpenStudy (vuriffy):

36 + 4x^2 = 4x^2 + 8x + 4

OpenStudy (vuriffy):

Solve for x, I would believe.

OpenStudy (jiteshmeghwal9):

Just replace 7 by 6 in my equation and solve for x

OpenStudy (vuriffy):

I got x = 4. 36 + 64 = 100. Lengths are 6, 8, and 10.

OpenStudy (triciaal):